OpenStudy (anonymous):

Can some one help me with my unit assessment from k12 its unit:3 lesson:13 its a math unit assessment

OpenStudy (anonymous):

I really could use help :D I will fan and medal! its for 8th grade

OpenStudy (anonymous):

pellet i am from k12 to um i am in 7th but i am in advanced tops what are you

OpenStudy (anonymous):

You can help me? XD im usually a good A student but this is confusing

OpenStudy (anonymous):

are you in tops or txva

OpenStudy (anonymous):

tops

OpenStudy (anonymous):

Which points have opposite x- and y-coordinates [for example, (2, –2) or (–3, 3)]? (Hint: Some answers may have some, but not all points that name opposites. Choose the answer that contains only points that have opposite x- and y-coordinates for each point named.) A. Z, L B. D, L C. D, Z D. C, D, Z

OpenStudy (anonymous):

lol me to um did you now that you are in grade 8th but you are doing 9th grade stuff so i am in 7th and i am doin 8th grade stuff XD LOL

OpenStudy (anonymous):

yes I new that, its pretty advanced XD

OpenStudy (anonymous):

yay ikr

OpenStudy (anonymous):

So can you help me...or...no?

OpenStudy (anonymous):

let me see hmmmmmmmmmmm!!!!!!!!!!!

OpenStudy (anonymous):

I'll fan and medal lol

OpenStudy (anonymous):

Which points have opposite x- and y-coordinates [for example, (2, –2) or (–3, 3)]? (Hint: Some answers may have some, but not all points that name opposites. Choose the answer that contains only points that have opposite x- and y-coordinates for each point named.) A. Z, L B. D, L C. D, Z D. C, D, Z

OpenStudy (anonymous):

almost got it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

slow wifi nooooooooooooooooooooooooooo!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

O.O no help? NOOOOOOOOOOO!!!!!!!!!

OpenStudy (anonymous):

no tat ok i will try my best

OpenStudy (anonymous):

Which points have opposite x- and y-coordinates [for example, (2, –2) or (–3, 3)]? (Hint: Some answers may have some, but not all points that name opposites. Choose the answer that contains only points that have opposite x- and y-coordinates for each point named.) A. Z, L B. D, L C. D, Z D. C, D, Z

OpenStudy (anonymous):

I think the answer is B. D,L am I right?

OpenStudy (anonymous):

https://static.k12.com/calms_media/media/308500_309000/308802/1/b69464e07231684ed1e041ba0948e7e92cbb5e68/MS_ALG_S1_07_02_quiz_Q3_question.gif Which point is the graph of the ordered pair (2, –4)? A. point A B. point B C. point C D. point D

OpenStudy (vheah):

You just have to look for OPPOSITE X AND Y COORDINATES. meaning that the two numbers inside the parentheses should be positive and negative but the same number. To do that you have to look at the graph, look at each letter and it's dot, and align it with the number line. the numbers on the horizontal line(the line that goes left to right) would be your x-coordinate and the numbers on your vertical line (the line that goes up and down) would hold your y-coordinates

OpenStudy (anonymous):

sorry lol i found the answer, but I need help on this new one, would you mind helping?

OpenStudy (anonymous):

https://static.k12.com/calms_media/media/308500_309000/308802/1/b69464e07231684ed1e041ba0948e7e92cbb5e68/MS_ALG_S1_07_02_quiz_Q3_question.gif Which point is the graph of the ordered pair (2, –4)? A. point A B. point B C. point C D. point D

OpenStudy (vheah):

Look at each letter and find the one that has x-coordinate 2 and y-coordinate -4

OpenStudy (anonymous):

B. point B?

OpenStudy (vheah):

Yep! :)

OpenStudy (anonymous):

thanks, I have a few more, I will fan and medal

OpenStudy (vheah):

Okay go ahead.

OpenStudy (anonymous):

In which quadrant does the point (10, –9) lie? A. Quadrant I B. Quadrant III C. Quadrant II D. Quadrant IV There is no graph...no picture either...

OpenStudy (anonymous):

how do I find this out, you dont have to tell me the answer, just explain how to do it if you'd like :D

OpenStudy (anonymous):

I think its B. Quadrant III am i right?

OpenStudy (anonymous):

@vheah ?

OpenStudy (vheah):

It would be quadrant IV. You have a positive 10 x-coordinate and a -9 y-coordinate

OpenStudy (anonymous):

oh i thought it was -10 and -9 oops XD

OpenStudy (anonymous):

can i ask another or...?

OpenStudy (vheah):

It's okay! Just remember that Quadrant III is all negative on the x-axis and the y-axis :)

OpenStudy (anonymous):

you are really helpful your awsome thanks so much

OpenStudy (vheah):

I'm here to help! :) So please go ahead. I still have some time

OpenStudy (anonymous):

YAY! thanks

OpenStudy (anonymous):

Which answer shows this equation written in standard form? 8 – 1(x – y) = –3x + 5 A. 2x + y = –3 B. 2x – y = –3 C. –4x + y = –3 D. 2x + y = 13 How do you do standard form

OpenStudy (anonymous):

I forgot...

OpenStudy (anonymous):

why am i getting tons of notifications about this q huh????????????????????????

OpenStudy (anonymous):

Which answer shows this equation written in standard form? 8 – 1(x – y) = –3x + 5 A. 2x + y = –3 B. 2x – y = –3 C. –4x + y = –3 D. 2x + y = 13 How do you do standard form again....?

OpenStudy (anonymous):

I think its A. 2x + y = –3

OpenStudy (anonymous):

@vheah ? I dont know if your there XD but i'll wait XD

OpenStudy (vheah):

That's right. sorry about that this thing is glitching, I'm in a public library rn! but that's right

OpenStudy (anonymous):

Oh, lol thanks agian, I have like...6 more questions...sadly if I dont pass, I will have to retake the whole thing again

OpenStudy (vheah):

A standard form would be ax+by=c you just have to get an answer that follows that structure but make sure you solved it right

OpenStudy (anonymous):

I have 6 more questions...can you help, im not trying to bother you with this though...

OpenStudy (vheah):

Alright. When is this thing due? I have a class 5:30 sadly, but I can do one more question. Hopefully I can answer it!

OpenStudy (anonymous):

Which set of points lies on the given graph? A. (5, 13), (1, 1), (2, –1) B. (–5, 13), (1, 1), (2, 1) C. (5, 13), (–1, 1), (2, 1) D. (–5, 13), (1, 1), (2, –1)

OpenStudy (vheah):

Okay, so the easiest thing to do here is since you're given a multiple choice, you can look at the line and see if it hits any of the ordered pairs on each choice.

OpenStudy (anonymous):

its not multiple choice

OpenStudy (anonymous):

so I think the answer is D. (–5, 13), (1, 1), (2, –1)

OpenStudy (vheah):

Im talking about the A, B, C, and D choices for this graph and the ordered pairs it gives you. It is not A and C because the x-coordinate does not go past positive an x-coordinate 5 while having a y-coordinate positive 13.

OpenStudy (vheah):

let me check

OpenStudy (anonymous):

is the answer D. (–5, 13), (1, 1), (2, –1)?

OpenStudy (vheah):

Yep!

OpenStudy (anonymous):

thanks, can anyone help me still though? I have 5 more questions...

OpenStudy (vheah):

Give me one more and ill try to help you with that one afterwards I have to get to class xD

OpenStudy (anonymous):

I dont want to seam like im just getting answers but I really need help

OpenStudy (vheah):

it's okay! You're figuring it all out so it's very good. I have 10 more minutes left before I have to get to class. If I fail to answer ur question now, I apologize in advance and I hope someone helps out!

OpenStudy (anonymous):

What is the equation of the line? https://static.k12.com/calms_media/media/1503500_1504000/1503772/1/5afd0b9bb4861bae518958d61ebaa2a6f9dddcc4/MS_IMC-140523-131303.jpg A. y=-2x B. y=-1/2x C. y=1/2x D. y=2x

OpenStudy (anonymous):

@vheah

OpenStudy (vheah):

This one is rise/run, see how many times it rises vertically and how many times it runs horizontally back to the line

OpenStudy (anonymous):

I believe its y=2x is that right?

OpenStudy (vheah):

Yep! :)

OpenStudy (anonymous):

thanks, you have time for one more or no?

OpenStudy (vheah):

I only have 5 minutes and ill be heading to class. I'm sorry! I hope someone helps you out and that I helped you some!

OpenStudy (anonymous):

ok thanks for all your help, im going to ask the other questions in another post, thanks for all your help!

OpenStudy (anonymous):

XD this question will now be closed

OpenStudy (vheah):

You're very welcome <3

OpenStudy (anonymous):

:)